### v1

@@ 4.1.1 RQ1: Which VR projects are subjected to heavy cloning?
We begin by examining the impact of different detection methods on the results, revealing a clear and consistent trend in the ${\sf NCF}$ metric (i.e., the number of clone functions): ${\sf NCF}$(Type 3-2) > ${\sf NCF}$(Type 3-2c) > ${\sf NCF}$(Type 3-1), and ${\sf NCF}$(Type 2) > ${\sf NCF}$(Type 2c) > ${\sf NCF}$(Type 1). This suggests that the blind renaming method consistently identifies the most clone fragments, with the consistent renaming method coming in second. This finding exactly aligns with the theoretical expectations outlined in NiCad [15], where blind renaming replaces all identifiers with a generic placeholder “$X_{n}$”, while consistent renaming assigns sequential identifiers “$X_{n}$”($n$ is a sequence number). The rationale is straightforward: if a code fragment is recognized as similar with identifiers renamed as “$X_{n}$”, it will also be recognized as similar when all identifiers are replaced with “$X$”. Additionally, we observe that ${\sf NCF}$(Type 3-x) > ${\sf NCF}$(Type x) for x = 1, 2, 2c. This outcome is intuitive since a fragment meeting a 0% dissimilarity threshold will inherently meet a 30% threshold. Based on these observations, Type 3-2 emerges as the most effective detection method, with its results encompassing those of other types. In terms of the ${\sf NCC}$ metric, there appears to be no clear correlation between the detection types and this metric. As a result, we select Type 3-2 as the standard for code clone detection in the following analyses.
@@
PID Project Name Language Type1 Type2 Type2c Type3-1 Type3-2 Type3-2c $|F|$
@@
1 BlenderXR C,Python 358 175 1578 601 1486 567 1731 716 4448 1226 3112 889 27822
2 gpac C 6 3 1421 259 1268 220 1195 312 2699 326 2089 278 10170
3 The-Seed-Link-Future C# 44 19 231 84 223 85 201 86 462 166 295 112 8475
4 open-brush C# 38 17 132 60 127 58 183 81 409 163 218 89 7548
5 UnityOculusAndroidVRBrowser C# 36 15 156 53 150 55 107 47 255 86 177 65 3716
6 UnityPlugin C# 12 6 53 21 52 21 60 24 113 43 85 33 2799
7 lovr C 2 1 84 33 77 30 77 31 232 64 148 53 2173
8 UnityGameTemplate C# 20 10 60 25 59 25 53 22 142 47 98 35 1981
9 ViveInputUtility-Unity C# 20 10 62 27 62 27 69 32 125 49 84 36 1863
10 com.xrtk.core C# 3 1 49 18 49 18 42 17 148 51 82 32 1607
@@ 9 ViveInputUtility-Unity C# 20 10 62 27 62 27 69 32 125 49 84 36 1863
11 SoundSpace C# 2149 1063 2190 1011 2185 1009 2175 985 2225 965 2201 1001 9117
12 RhythmAttack-VR C# 37 17 162 60 149 62 118 48 330 119 221 89 6751
13 Dungeon-VR C# 20 10 100 38 91 35 103 41 254 98 154 61 6470
14 Situated-Empathy-in-VR C# 23 11 140 59 136 58 136 59 388 140 262 100 4788
15 mineRVa C# 18 9 88 35 82 34 53 24 197 71 140 58 4722
16 VR-Escape-Room C# 12 6 53 19 44 16 31 13 120 46 77 30 4067
17 Group6_ProjectNurture C# 27 10 148 49 142 51 102 44 248 82 171 62 3980
18 vrtist C# 51 17 114 42 112 41 141 48 299 103 187 69 3322
19 Terminal C# 19 8 96 27 88 29 71 27 160 49 121 41 2929
20 elite-vr-cockpit C# 0 0 52 18 49 17 30 14 109 39 68 26 2725
To expand our study on cloning in serialized asset files within VR software, we add 10 more VR applications, covering diverse subcategories like games, education, simulators, and AR, ensuring a comprehensive analysis. Development tools are excluded from the analysis of file cloning because they are typically source-code-centric, with asset files serving configuration and descriptive purposes. In contrast, asset files dominate in application projects, with source code primarily existing as scripts to manipulate these assets. Table 5 provides the clone detection results for the 20 selected VR software applications. We observe that the three projects with the highest number of file clones (${\sf NCA}$)—“Dungeon-VR”, “Group6_ProjectNurture” and “OpendagVR2”—are also the ones with the largest number of asset files ($|A|$), following a pattern similar to source code cloning.
@@ 19 Terminal C# 19 8 96 27 88 29 71 27 160 49 121 41 2929
11 SoundSpace 207 10 15 3 28 391 52.94% 7.16% 11.24%
12 RhythmAttack-VR 286 20 19 5 44 658 43.47% 6.69% 8.17%
13 Dungeon-VR 941 31 20 15 66 1737 54.17% 3.80% 16.32%
14 Situated-Empathy-in-VR 158 1 3 2 6 209 75.60% 2.87% 57.42%
15 mineRVa 235 8 8 3 19 411 57.18% 4.62% 18.67%
16 VR-Escape-Room 154 10 6 4 20 355 43.38% 5.63% 9.69%
17 Group6_ProjectNurture 777 24 15 12 51 1099 70.70% 4.64% 23.28%
18 vrtist 293 21 17 5 43 408 71.81% 10.54% 14.89%
19 Terminal 268 11 13 4 28 450 59.56% 6.22% 23.18%
20 elite-vr-cockpit 97 5 8 1 14 212 45.75% 6.60% 12.35%
21 VRHamsterBall 108 5 10 1 16 194 55.67% 8.25% 16.37%
22 OpendagVR2 699 21 26 10 57 947 73.81% 6.02% 24.05%
23 epicslash 145 7 8 6 21 210 69.05% 10.00% 17.46%
24 CityMatrixAR 70 3 5 2 10 115 60.87% 8.70% 25.13%
25 Procrastination-VR 438 5 5 2 12 478 91.63% 2.51% 69.29%
26 AirAttack 265 8 9 5 22 352 75.28% 6.25% 27.28%
27 GolfVR 56 4 3 2 9 116 48.28% 7.76% 13.84%
28 XR-Keyboard 70 5 2 2 9 159 44.03% 5.66% 15.18%
29 HorrorGame 349 12 27 7 46 506 68.97% 9.09% 13.93%
30 Pokemon-Themed-Kiosk-VR 172 10 4 5 19 239 71.97% 7.95% 22.96%
@@ 4.1.2 RQ2: What are the main differences in code cloning between VR and non-VR software?
To answer this RQ, we suggest exploring the relationship between the quantity of asset files ($|A|$) and the count of clone functions (${\sf NCF}$), as VR software differs from traditional software primarily in its usage of serialized asset files.
As shown in Table 6, the number of cloned functions and serialized asset files for each of the 30 projects is presented. We observe that the development tools Project-1 (“BlenderXR”), Project-2 (“gpac”), and the VR applications Project-11 (“SoundSpace”), Project-24 (“CityMatrixAR”) exhibit significantly higher levels of source code cloning than the rest of the projects. Considering that Project-1 and Project-2 neither use C# nor contain serialized asset files, these two projects can effectively be viewed as resembling non-VR software.
Table 6: The number of asset files and clone functions in the projects.
@@
Among the remaining eight development tools projects, seven are Unity-based, which means they make use of considerable reusable code stored in serialized files. This results in a lower level of code cloning, as tools like NiCad are less effective at identifying clones within serialized files, which differ from standard source code. Project-7 (“lovr”), although written in C, also contains many model class files (/modules), which contributes to its relatively lower levels of code cloning. This observation does not contradict our conclusions.
In terms of Project-11 and Project-24, we find that they have relatively few serialized files in relation to their source code volume, causing higher levels of code cloning. In fact, Project-24, being an AR project, constructs models from real-world data rather than serialized files. Similarly, Project-11, an acoustic simulation and visualization project, does not require extensive use of serialized files. The remaining 18 projects maintain a relatively low level of code cloning.
In summary, the code cloning levels of VR software are generally lower than those of traditional software. This is because processes that tend to generate clones, such as class and method creation, are often implicitly stored within asset files. As a result, the cloning metrics might misleadingly appear lower. The impact of asset files and their relationship to cloning cannot be ignored, adding further challenges for clone detection in VR software.
Insight 2.  The main difference between code clones in VR software and regular software lies in the presence of serialized files, which significantly reduce the level of source code cloning in a project.
4.1.3 RQ3: How to identify the most appropriate metrics for VR cloning analysis?
@@
Insight 3.  The proposed metrics measure cloning from distinct perspectives: the number or ratio of clone functions (files) measures the quantity of clones, while the number or ratio of clone classes (groups) evaluates clone concentration. By combining these metrics, a comprehensive clone evaluation can be achieved.
@@ 4.2 Source Code Cloning Analysis
To further examine which programming languages are more prone to code cloning, we identify projects with 10 or more clone classes ($\sf NCC$>= 10) in terms of Clone Type 3-2 within high-star VR projects and categorize these projects by their primary programming languages. Table 8 presents the distribution of projects across various languages at different cloning levels. The results reveal that C# accounts for the highest number of projects (18) with over 10 clone classes, followed by C with 5 projects. This finding aligns with the prominent role of C# as the most commonly used language in VR software development. However, when considering the cloning ratio, C ($5/11\approx 45.5\%$) surpasses C# ($18/46\approx 39.1\%$), indicating that C may have a higher likelihood of triggering cloning issues. This may stem from the fact that C, being process-oriented, does not support the code reuse mechanisms inherent to object-oriented languages, leading to more code cloning. In contrast, the presence of serialized files in C# programs leads to the occurrence of code clones at lower levels, as demonstrated in RQ2.
Insight 4.  C# is the most commonly used programming language in open-source VR projects, whereas C is more likely to cause source code cloning issues. However, C# introduces an additional challenge with the cloning of serialized files.
@@ 4.2.2 RQ5: How do third-party libraries impact code cloning in VR software?
To answer this RQ, we search through the 20 application projects to identify the presence of widely used third-party libraries (e.g., SteamVR, Oculus, OpenVR, Vive, Google VR, VRTK). The results are summarized in Table 9, listing only the projects with code clones induced by third-party libraries, along with the specific libraries involved.
Upon reviewing our results, we find that despite exploring various well-known VR third-party libraries, code clones are only successfully identified in projects using SteamVR, Oculus, and VRTK. Among these, SteamVR appears in 10 projects, while Oculus and VRTK each appear in 4 projects. However, when it comes to introducing source code clones, SteamVR has the lowest contribution, Oculus ranks in the middle, and VRTK contributes the most. This is reasonable because VRTK, as an open-source toolkit, supports numerous VR devices and platforms, providing high-level abstractions that simplify development and lead to significant code redundancy. On the other hand, SteamVR and Oculus are development kits designed for specific hardware, prioritizing compatibility and low-level optimization, which results in less cross-platform or high-level redundant code.
@@
Insight 5.  Third-party libraries introduced in VR applications often contribute to source code cloning issues, sometimes constituting the majority of the cloning results. The more functionalities the third-party libraries provide, the higher the clone evaluation metrics tend to be.
4.2.3 RQ6: How do inter-version and intra-version code cloning evolve across different versions of a VR project?
@@
Insight 6.   Inter-version code cloning is more common than intra-version code cloning, introducing significant security risks due to the potential propagation of vulnerabilities across multiple versions of the software.

### v2

@@ 4.1.1 RQ1: Which VR projects are subjected to heavy cloning?
We begin by examining the impact of different detection methods on the results, revealing a clear and consistent trend in the $\sf NCF$} metric (i.e., the number of clone functions): $\sf NCF$}(Type 3-2) > $\sf NCF$}(Type 3-2c) > $\sf NCF$}(Type 3-1), and $\sf NCF$}(Type 2) > $\sf NCF$}(Type 2c) > $\sf NCF$}(Type 1). This suggests that the blind renaming method consistently identifies the most clone fragments, with the consistent renaming method coming in second. This finding exactly aligns with the theoretical expectations outlined in NiCad [15], where blind renaming replaces all identifiers with a generic placeholder $X_{n}$”, while consistent renaming assigns sequential identifiers $X_{n}$”($n$ is a sequence number). The rationale is straightforward: if a code fragment is recognized as similar with identifiers renamed as $X_{n}$”, it will also be recognized as similar when all identifiers are replaced with “XX”. Additionally, we observe that $\sf NCF$}(Type 3-x) > $\sf NCF$}(Type x) for x = 1, 2, 2c. This outcome is intuitive since a fragment meeting a 0% dissimilarity threshold will inherently meet a 30% threshold. Based on these observations, Type 3-2 emerges as the most effective detection method, with its results encompassing those of other types. In terms of the $\sf NCC$} metric, there appears to be no clear correlation between the detection types and this metric. As a result, we select Type 3-2 as the standard for code clone detection in the following analyses.
@@
PID Project Name Domain Language Type1 Type2 Type2c Type3-1 Type3-2 Type3-2c $|F|$
@@
1 BlenderXR / C,Python 358 175 1578 601 1486 567 1731 716 4448 1226 3112 889 27822
2 gpac / C 6 3 1421 259 1268 220 1195 312 2699 326 2089 278 10170
3 The-Seed-Link-Future / C# 44 19 231 84 223 85 201 86 462 166 295 112 8475
4 open-brush / C# 38 17 132 60 127 58 183 81 409 163 218 89 7548
5 UnityOculusAndroidVRBrowser / C# 36 15 156 53 150 55 107 47 255 86 177 65 3716
6 UnityPlugin / C# 12 6 53 21 52 21 60 24 113 43 85 33 2799
7 lovr / C 2 1 84 33 77 30 77 31 232 64 148 53 2173
8 UnityGameTemplate / C# 20 10 60 25 59 25 53 22 142 47 98 35 1981
9 ViveInputUtility-Unity / C# 20 10 62 27 62 27 69 32 125 49 84 36 1863
10 com.xrtk.core / C# 3 1 49 18 49 18 42 17 148 51 82 32 1607
@@ 9 ViveInputUtility-Unity / C# 20 10 62 27 62 27 69 32 125 49 84 36 1863
11 SoundSpace simulator C# 2149 1063 2190 1011 2185 1009 2175 985 2225 965 2201 1001 9117
12 RhythmAttack-VR gaming C# 37 17 162 60 149 62 118 48 330 119 221 89 6751
13 Dungeon-VR gaming C# 20 10 100 38 91 35 103 41 254 98 154 61 6470
14 Situated-Empathy-in-VR education C# 23 11 140 59 136 58 136 59 388 140 262 100 4788
15 mineRVa education C# 18 9 88 35 82 34 53 24 197 71 140 58 4722
16 VR-Escape-Room gaming C# 12 6 53 19 44 16 31 13 120 46 77 30 4067
17 Group6_ProjectNurture healthcare C# 27 10 148 49 142 51 102 44 248 82 171 62 3980
18 vrtist education C# 51 17 114 42 112 41 141 48 299 103 187 69 3322
19 Terminal gaming C# 19 8 96 27 88 29 71 27 160 49 121 41 2929
20 elite-vr-cockpit gaming C# 0 0 52 18 49 17 30 14 109 39 68 26 2725
To expand our study on cloning in serialized asset files within VR software, we add 10 more VR applications, covering diverse subcategories like games, education, simulators, and AR, ensuring a comprehensive analysis. Development tools are excluded from the analysis of file cloning because they are typically source-code-centric, with asset files serving configuration and descriptive purposes. In contrast, asset files dominate in application projects, with source code primarily existing as scripts to manipulate these assets. Table 7 provides the clone detection results for the 20 selected VR software applications.
We observe that the three projects with the highest $\sf NCA$} (i.e., the number of file clones)—“Dungeon-VR”, “Group6_ProjectNurture” and “OpendagVR2”—are also the ones with the largest $|A|$ (i.e., the number of asset files), following a pattern similar to source code cloning. For better clarity, Figure 2 visually depicts the relationship between project characteristics and clone metrics, specifically the correlation between the number of asset files ($|A|$) and file clones ($\sf NCA$}), with a near-linear relationship observed, suggesting that larger projects typically exhibit more file clones.
Figure 2: The relationship between $|A|$ and $\sf NCA$}.
@@ 19 Terminal gaming C# 19 8 96 27 88 29 71 27 160 49 121 41 2929
12 RhythmAttack-VR gaming 286 20 19 5 44 658 43.47% 6.69% 8.17%
13 Dungeon-VR gaming 941 31 20 15 66 1737 54.17% 3.80% 16.32%
16 VR-Escape-Room gaming 154 10 6 4 20 355 43.38% 5.63% 9.69%
19 Terminal gaming 268 11 13 4 28 450 59.56% 6.22% 23.18%
20 elite-vr-cockpit gaming 97 5 8 1 14 212 45.75% 6.60% 12.35%
21 VRHamsterBall gaming 108 5 10 1 16 194 55.67% 8.25% 16.37%
23 epicslash gaming 145 7 8 6 21 210 69.05% 10.00% 17.46%
26 AirAttack gaming 265 8 9 5 22 352 75.28% 6.25% 27.28%
27 GolfVR gaming 56 4 3 2 9 116 48.28% 7.76% 13.84%
28 XR-Keyboard gaming 70 5 2 2 9 159 44.03% 5.66% 15.18%
29 HorrorGame gaming 349 12 27 7 46 506 68.97% 9.09% 13.93%
30 Pokemon-Themed-Kiosk-VR gaming 172 10 4 5 19 239 71.97% 7.95% 22.96%
14 Situated-Empathy-in-VR education 158 1 3 2 6 209 75.60% 2.87% 57.42%
15 mineRVa education 235 8 8 3 19 411 57.18% 4.62% 18.67%
18 vrtist education 293 21 17 5 43 408 71.81% 10.54% 14.89%
22 OpendagVR2 education 699 21 26 10 57 947 73.81% 6.02% 24.05%
25 Procrastination-VR education 438 5 5 2 12 478 91.63% 2.51% 69.29%
17 Group6_ProjectNurture healthcare 777 24 15 12 51 1099 70.70% 4.64% 23.28%
11 SoundSpace simulator 207 10 15 3 28 391 52.94% 7.16% 11.24%
24 CityMatrixAR simulator 70 3 5 2 10 115 60.87% 8.70% 25.13%
@@
When examining the domains of the top 20 most clone-prone software systems, gaming and education emerge as the most impacted. According to Table 6 and Table 7, code cloning occurs in 5 gaming and 3 education systems, while file cloning is present in 12 gaming and 5 education systems. This recurring pattern highlights the need for special attention to cloning control in these areas, especially in gaming.
Insight 2.Code and file cloning issues are most prevalent in gaming and education software, with gaming being the most significantly impacted, highlighting the need for targeted attention in this domain.
@@ 4.1.2 RQ2: What are the main differences in code cloning between VR and non-VR software?
To answer this RQ, we start by reviewing prior work on code cloning in traditional software. Among the most relevant studies, Kamiya et al. [13] reported a cloned function ratio of 12% to 16% in medium and large software. Kapser et al. [72] found a similar ratio, close to 20%, in C-based projects. Roy et al. [71] observed that Java and C++ projects exhibited a clone function ratio ranging from 10% to 15%. These findings suggest that code cloning is a widespread and consistent issue across traditional software development, typically affecting 10% to 20% of functions.
To investigate code cloning in VR software, we propose exploring the relationship between the ratio of clone functions ($\sf rcf$}) and the quantity of asset files ($|A|$), as the use of such assets constitutes a key difference between VR and traditional software. As shown in Table 8, we observe that the development tools Project-1 (“BlenderXR”) and Project-2 (“gpac”), along with the VR applications Project-11 (“SoundSpace”) and Project-24 (“CityMatrixAR”) exhibit significantly higher levels of source code cloning than other projects, with $\sf rcf$} values ranging from 16% to 27%, whereas most of the remaining 26 projects fall below 10%.
We further examine the four outliers with unusually high levels of code cloning. Project-1 and Project-2 lack any serialized asset files. Project-11 and Project-24, though containing some serialized files, have relatively few when compared to the overall volume of source code. Specifically, Project-24, an AR application, constructs models from real-world inputs rather than serialized data. Similarly, Project-11, which focuses on acoustic simulation and visualization, does not require extensive use of serialized assets. A shared trait among these projects is the limited presence of serialized files, which aligns them more closely with traditional software patterns. As a result, they exhibit cloning behavior similar to that observed in traditional software, with clone function ratios exceeding 10%.
By comparison, VR software projects that utilize extensive serialized asset files demonstrate markedly reduced code cloning, as evidenced by the remaining 26 projects’ average clone function ratio of 5.02%. This may result from the limited capability of clone detection tools (e.g., NiCad) to recognize duplicates in serialized files, which differ structurally from standard source code. It is important to note that Project-7 (“lovr”), although written in C, also contains many model class files (/modules), which contributes to its relatively lower levels of code cloning. This observation does not contradict our conclusions. In summary, asset files significantly influence code cloning patterns, presenting unique detection challenges that current clone analysis tools must address for effective VR software maintenance.
Table 8: The correlation between the ratio of clone functions ($\sf rcf$}) and the number of asset files ($|A|$) across VR projects.
@@
Insight 3.Unlike traditional software, most VR projects exhibit significantly lower source code cloning due to their heavy reliance on serialized asset files.
4.1.3 RQ3: How to measure and analyze code cloning in VR software development?
@@
Insight 4.The proposed metrics measure cloning from distinct perspectives: the number or ratio of clone functions (files) measures the quantity of clones, while the number or ratio of clone classes (groups) evaluates clone concentration. By combining these metrics, a comprehensive clone evaluation can be achieved.
@@ 4.2 Source Code Cloning Analysis
To further examine which programming languages are more prone to code cloning, we identify projects with 10 or more clone classes ($\sf NCC$>= 10) in terms of Clone Type 3-2 within high-star VR projects and categorize these projects by their primary programming languages. Table 10 presents the distribution of projects across various languages at different cloning levels. The results reveal that C# accounts for the highest number of projects (18) with over 10 clone classes, followed by C with 5 projects. This finding aligns with the prominent role of C# as the most commonly used language in VR software development. However, when considering the cloning ratio, C ($5/11\approx 45.5\%$) surpasses C# ($18/46\approx 39.1\%$), indicating that C may have a higher likelihood of triggering cloning issues. This may stem from the fact that C, being process-oriented, does not support the code reuse mechanisms inherent to object-oriented languages, so developers might inadvertently copy blocks of code instead of abstracting them.
Insight 5.While C# is the most commonly used programming language in open-source VR projects, C-based projects exhibit a higher proportion of code cloning.
@@ 4.2.2 RQ5: How do third-party libraries impact code cloning in VR software?
Table 11: The list of analyzed third-party libraries.
Library Type Key Files Exist Clone
SteamVR VR Hardware SDK Assets/SteamVR/Input/SteamVR_Input.cs
Oculus VR Platform SDK Assets/Oculus/VR/Scripts/OVRManager.cs
OpenVR VR Runtime Assets/Plugins/openvr/api.cs ×
Vive Hardware Interaction SDK Assets/SteamVR/InteractionSystem/Core/Interfaces/Hand.cs ×
Google VR Mobile VR SDK Assets/GoogleVR/Legacy/Scripts/GvrViewer.cs ×
VRTK VR Framework Assets/VRTK/SDK/Base/Scripts/VRTK_SDKManager.cs
Unity XR Cross-Platform SDK Assets/XR/Input/XRController.cs ×
OpenXR Cross-Vendor Runtime Assets/XR/OpenXR/Settings/OpenXRSettings.asset ×
OSVR VR Platform Assets/OSVR/ClientKit/OSVR_ClientContext.cs ×
Pico SDK VR SDK Assets/PicoMobileSDK/Scripts/PicoVR_Manager.cs ×
Wave SDK VR Platform Assets/WaveVR/Scripts/WaveVR_Reticle.cs ×
Varjo Base Enterprise VR SDK Assets/Varjo/Scripts/VarjoManager.cs ×
Ultraleap Gesture SDK Assets/Ultraleap/Hands/LeapHandController.cs ×
WebXR Browser-Based VR Assets/WebXR/Plugins/WebXRInterface.cs ×
@@
To answer this RQ, we examine 20 application projects to identify third-party libraries with code clones. We initially identify 14 widely used libraries, 6 of which are found in our dataset, with 3—SteamVR, Oculus, and VRTK—showing cloning issues. The specific procedure is as follows: (i) we identify 14 widely used third-party libraries in VR development for study, as shown in Table 11; (ii) we verify whether each of these 14 libraries is integrated in any of the 20 projects by checking for the existence of their distinctive key files, resulting in 6 libraries being present; (iii) we then detect whether the remaining libraries have code clones, identifying 3 libraries—SteamVR, Oculus, and VRTK—with cloning issues; and (iv) we review the cloning results to ensure all libraries with clones are properly examined.
Upon reviewing our results, we find that SteamVR appears in 10 projects, while Oculus and VRTK each appear in 4 projects. However, when it comes to introducing source code clones, SteamVR has the lowest contribution, Oculus ranks in the middle, and VRTK contributes the most. This is reasonable because VRTK, as an open-source toolkit, supports numerous VR devices and platforms, providing high-level abstractions that simplify development and lead to significant code redundancy. On the other hand, SteamVR and Oculus are development kits designed for specific hardware, prioritizing compatibility and low-level optimization, which results in less cross-platform or high-level redundant code.
@@
Insight 6.Third-party libraries introduced in VR applications often contribute to source code cloning issues, sometimes constituting the majority of the cloning results. The more functionalities the third-party libraries provide, the higher the clone evaluation metrics tend to be.
4.2.3 RQ6: How do intra-version and inter-version code cloning evolve across different versions of a VR project?
@@
Insight 7. Inter-version code cloning is more common than intra-version code cloning, introducing significant security risks due to the potential propagation of vulnerabilities across multiple versions of the software.